# DashCLIP: Leveraging multimodal models for generating semantic embeddings for DoorDash

### v1

@@ 4.2.1. A/B Testing & Rollout
To evaluate the best model candidate in production, we created an online A/B experiment to serve real users on the DoorDash platform. This consumer-bucketed test aimed to measure the business metric improvement of the new treatment model over the control DCN baseline. It was held in August 2024 for 10 days on the in-store surface. The experiment analysis showed that the new model increased the engagement rates for most of the top queries and categories, driving more revenue for the Sponsored Products advertisement and improving the relevance measure. The quantitative gain is reported in Table 2. With the positive result of top-line metrics, the model is fully deployed to serve 100% of the traffic by the end of the same month.
@@ 4.2.2. Online Serving
In order to retrieve the pre-trained embedding for online inference, we compared two viable strategies: feature fetching and model fusion. For the former, we fetch the embedding features by doing product / user / query id lookups from the online feature store and feed the embeddings to the model. For the latter, we implant embedding tables directly in the model definition by extending the EmbeddingBagCollection module of the torchrec library to fuse the tables with the model. These strategies present a trade-off between network overhead induced by fat feature fetches of the actual embedding values vs memory constraint of the larger fused model size. Eventually we serve with the latter approach because latency will increase substantially due to fetches when the data and model are not co-located closely. However, if the constraint is on the GPU’s memory, it will be better not to fuse the model and embedding tables.
@@ 5. E-commerce Applications
Since our goal is to develop generalized product and query embeddings for different applications, we explore DashCLIP in tasks beyond search ranking. To this, we pick aisle category prediction and product-query relevance prediction as our two additional tasks. These tasks are important to improve DoorDash’s overall user experience and analogous tasks are prevalent across other e-commerce platforms. Accurate product categorization ensures seamless product discovery, and query-based relevance filtering is crucial to improve the user search experience. Together, these capabilities would improve the overall engagement and browsing experience on any e-commerce platform. Further, they test different aspects of our framework against the off-the-shelf BLIP model: the aisle-category prediction checks whether the final embeddings are well-adapted to our domain, while the product-query relevance task accounts for the alignment between Product and Query encoders. We perform both qualitative and quantitative evaluations using set-aside 2% of the Query-Product Relevance Dataset with stratification by aisle category. The details of the tasks are as follows:
@@ 5.1. Aisle Category Prediction
We further perform an 80-20 stratified split on the unique product ids within the set-aside data to create the training and testing datasets. The dataset sizes are 13500 and 3375, respectively. For the baseline, we use the BLIP-14M checkpoint and compare it with DashCLIP. We train a simple classifier on top of the image-text Product Encoder by passing the last embedding output of the encoder through a hidden layer, followed by dropout and a final linear layer. We keep the encoder frozen and only train the added layers. The results are shown in Table 3. We observe that our model performs significantly better with an average F1 score of 0.850 compared to 0.801 for the baseline. The baseline model tends to perform poorer on recall for low-support classes, whereas our model is able to limit this issue. From the qualitative evaluation results shown in Figure 4, we observe that the classes are naturally clustered, and semantically similar clusters are closer to each other.
@@ 5.2. Product-Query Relevance Prediction
To prepare the dataset for this task, we first sample 10% of data points from the kept-aside dataset. Next, due to the class imbalance in the Query Product Relevance dataset, we randomly drop 50% of IR class samples to remove easy negatives. Finally, we perform an 80-20 split on the obtained data to generate the train and test sets. The final sizes are 32577 and 8175, respectively. We again use the BLIP-14M checkpoint as the baseline. To build the classifier, we extract the multi-modal product embedding from the image-text encoder and concatenate the query embedding to pass it through a dropout layer followed by a classification layer. For the BLIP-14M, we encode the queries using the text encoder, and for DashCLIP, the query encoder is utilized to generate the query embedding. Again, we only train the added layers. As reported in Table 4, our model achieves an average F1 score of 0.671 compared to 0.476 for the baseline. Similar to the first task, compared to our model, the baseline model achieves an extremely low recall for the HR class due to its low frequency in our data. From the qualitative evaluation shown in Figure 5, DashCLIP is able to create significantly more separation between the three classes compared to the baseline.
@@ 6. Conclusion
Best Model Baseline Model
Aisle Category Precision Recall F1 Precision Recall F1
Alcohol 0.973 0.972 0.973 0.966 0.944 0.955
Baby & Child* 0.849 0.816 0.832 0.932 0.724 0.815
Bakery* 0.861 0.633 0.729 0.875 0.429 0.575
Candy 0.882 0.882 0.882 0.821 0.836 0.829
Dairy & Eggs* 0.931 0.779 0.848 0.868 0.767 0.815
Drinks 0.907 0.947 0.927 0.863 0.954 0.906
Frozen 0.879 0.917 0.897 0.831 0.903 0.866
Household* 0.939 0.620 0.747 0.913 0.420 0.575
Meat & Fish* 0.824 0.737 0.778 0.867 0.684 0.765
Medicine 0.833 0.923 0.876 0.796 0.918 0.853
Pantry 0.769 0.792 0.780 0.746 0.762 0.754
Personal Care 0.940 0.926 0.933 0.908 0.932 0.919
Pet Care 0.997 0.995 0.996 0.987 0.995 0.991
Snacks 0.816 0.886 0.850 0.791 0.839 0.814
Vitamins* 0.938 0.566 0.706 0.957 0.415 0.579
Average 0.889 0.826 0.850 0.875 0.768 0.801
Table 3. Comparison of DashCLIP embedding with the baseline BLIP-14M for Aisle Category prediction task. * represents low-support classes (support ¡ 100 in the test set). We observe significant improvement in recall for these classes such as Bakery and Household.
Best Model Baseline Model
Relevance Label Precision Recall F1 Precision Recall F1
Irrelevant (IR) 0.805 0.900 0.850 0.693 0.891 0.78
Moderately Relevant (MR) 0.700 0.590 0.640 0.607 0.487 0.54
Highly Relevant (HR) 0.581 0.474 0.522 0.547 0.059 0.107
Average 0.695 0.655 0.671 0.616 0.479 0.476
Table 4. Comparison of DashCLIP with the baseline BLIP-14M for Product-Query Relevance prediction task. We achieve a higher performance in all three metrics, and a significant recall jump for HR class.
@@ 7.1. Pre-trained Embeddings
Contextualized representations, or embeddings, derived from pre-trained models have demonstrated state-of-the-art performance across various natural language (Devlin, 2018) and vision tasks (Dosovitskiy, 2020). Early work primarily focused on learning word-level representations. For instance, Word2Vec (Mikolov, 2013) introduced the continuous bag of words (CBOW) and Skip-gram models, which learn word representations based on neighboring context, while GloVe (Pennington et al., 2014) used word co-occurrence statistics to capture semantic relationships. ELMo (Sarzynska-Wawer et al., 2021), with its LSTM-based architecture, enabled context-dependent word representations, addressing some of the limitations of earlier models.
@@
The advent of Transformer architecture (Vaswani, 2017) revolutionized pre-trained language models, giving rise to widely-used models such as BERT (Devlin, 2018) and RoBERTa (Liu, 2019). Subsequent research extended these models to sentence-level embeddings (Reimers, 2019) (Gao et al., 2021), making them more suitable for tasks like semantic textual similarity (STS) (Cer et al., 2017). More recently, approaches have expanded to learn vision (Radford et al., 2021; Dosovitskiy, 2020) and multi-modal (Li et al., 2021, 2022, 2023) representations, enabling zero-shot transfer learning for vision-language tasks.
@@
Despite their success, pre-trained embeddings often lack the necessary precision for specific use cases, as they are typically trained on general datasets. Therefore, they need to be adapted or aligned to capture the particular entities and intents relevant to a specific platform. In e-commerce and social media applications, embeddings have been widely used for tasks such as personalized recommendations and search. For example, (Nigam et al., 2019) trained a deep neural network using customer activity logs to learn semantic representations of products. In the realm of multi-modal representations, MARN (Li et al., 2020) combined modality-specific and modality-invariant representations via an adversarial network, achieving state-of-the-art performance on several public datasets for CTR prediction. Similarly, (Baltescu et al., 2022) proposed a multi-task learning approach that unifies user, image, and search-based recommendations, utilizing different engagement signals. Additionally, (Xv et al., 2022) applied contrastive learning to generate visual representations for CTR prediction while addressing the performance degradation caused by selection sampling bias. Beyond product representations, numerous studies (Zhang et al., 2024; Pancha et al., 2022; Zhang et al., 2020) focus on learning user representations on online platforms.
7.2. CTR Prediction
@@
CTR prediction is crucial for e-commerce platforms and aims to predict the probability of a user clicking an impression. Earlier works relied on traditional machine learning techniques like logistic regression to model this probability using features such as CTR of related terms (Richardson et al., 2007) and number of impressions and card position (Kumar et al., 2015). Kernel SVM methods (Chang et al., 2010) were introduced to tackle limited feature interactions, while Factorization Machines (FMs) (Rendle, 2010; Juan et al., 2016; Pan et al., 2018) and ensemble-based gradient boosting models (Ling et al., 2017) aimed to address the data sparsity issues and improve feature selection, respectively. Deep Learning models lead to improved performance due to their ability to learn complex non-linear interactions: Product-based Neural Networks (PNN) (Qu et al., 2016) captures category interactions though embedding layers. The Wide and Deep architecture (Cheng et al., 2016) uses a combination of wide cross-product transformations for feature interactions memorization and deep neural network for generalization, training both components jointly. Making cross-feature selection more efficient, the Deep and Cross Network (DCN) (Wang et al., 2017) introduces the cross-network with learned explicit feature cross to reduce the need for manual feature engineering. DeepFM (Guo et al., 2017) unified low- and high- order feature interaction learning using a factorization machine (FM) based neural network. FiBiNET (Huang et al., 2019) uses a bilinear function to dynamically learn feature interactions and demonstrates SoTA performance combining with both shallow and deep models. Finally, DCNV2 (Wang et al., 2021) improves on the DCN model, allowing large-scale learning in industrial settings. While many of these models emphasize architectural enhancements, our approach remains architecture-agnostic, focusing on generating semantic embeddings that can be seamlessly integrated into any architecture.

### v2

@@ 4.3.1. A/B Testing & Rollout
To evaluate the best model candidate in production, we created an online A/B experiment to serve real users on our platform. This consumer-bucketed test aimed to measure the business metric improvement of the new treatment model over the control DCN baseline. It was held in August 2024 for 10 days on the in-store surface. The experiment analysis showed that the new model increased the engagement rates for most of the top queries and categories, driving more revenue for the Sponsored Products advertisement and improving the relevance measure. The quantitative gain is reported in Table 3. With the positive result of top-line metrics, the model is fully deployed to serve 100% of the traffic by the end of the same month. Compared to labeling and training cost reported in Section 3.2.3, the business revenue gain justified the large return on investment and deployment.
@@ 4.3.2. Online Serving
In order to retrieve the pre-trained embedding for online inference, we compared two viable strategies: feature fetching and model fusion. For the former, we fetch the embedding features by doing product / user / query id lookups from the online feature store and feed the embeddings to the model. For the latter, we implant embedding tables directly in the model definition by extending the EmbeddingBagCollection module of the torchrec library to fuse the tables with the model. These strategies present a trade-off between network overhead induced by fat feature fetches of the actual embedding values vs memory constraint of the larger fused model size. Eventually we serve with the latter approach because latency will increase substantially due to fetches when the data and model are not co-located closely. For instance, the model fusing approach increases the p99 latency by less than 4ms while feature fetching typically consumes 5x more latency budget than model inference.
@@ 5. E-commerce Applications
Given our goal to develop generalized embeddings for different applications, we explore DashCLIP in tasks beyond search recommendation. We pick aisle category prediction and product-query relevance prediction as two additional tasks. These tasks are important to improve the overall user experience on our platform and analogous tasks are prevalent across other e-commerce platforms: Accurate product categorization ensures seamless product discovery, and query-based relevance filtering is crucial to improve the user search experience. Further, the aisle-category prediction checks if the DashCLIP embeddings are aligned to our domain and product-query relevance tests the alignment between product and query encoders. Together, they perform a comprehensive evaluation of our model. We perform both qualitative and quantitative evaluations using the set-aside 2% of the Query-Product Relevance Dataset with stratification by aisle category. The details of the tasks are as follows:
@@ 5.1. Aisle Category Prediction
We further perform an 80-20 stratified split on the unique product ids within the set-aside data to create the training and testing datasets. The dataset sizes are 13500 and 3375, respectively. For the baseline, we use the BLIP-14M checkpoint and compare it with DashCLIP. We train a simple classifier on top of the image-text Product Encoder by passing the last embedding output of the encoder through a hidden layer, followed by dropout and a final linear layer. We keep the encoder frozen and only train the added layers.
We observe that our model performs significantly better with an average F1 score of 0.750 compared to 0.707 for the baseline. The baseline model tends to perform poorer on recall for low-support classes, whereas our model is able to limit this issue. From the qualitative evaluation results shown in Figure 4, we observe that the classes are naturally clustered, and semantically similar clusters are closer to each other.
@@ 5.2. Product-Query Relevance Prediction
To prepare the dataset for this task, we first sample 10% of data points from the set-aside dataset. Next, due to the class imbalance in the Query Product Relevance dataset from Section 2.2, we randomly drop 50% of IR class samples to remove easy negatives. Finally, we perform an 80-20 split on the obtained data to generate the train and test sets. The final sizes are 32577 and 8175, respectively. We again use the BLIP-14M checkpoint as the baseline. To build the classifier, we extract the multi-modal product embedding from the image-text encoder and concatenate the query embedding to pass it through a dropout layer followed by a classification layer. For the BLIP-14M, we encode the queries using the text encoder, and for DashCLIP, the query encoder is utilized to generate the query embedding. Again, we only train the added layers.
Our model achieves an average F1 score of 0.671 compared to 0.476 for the baseline. Similar to the first task, compared to our model, the baseline model achieves an extremely low recall (0.059) for the HR class due to its low frequency. From the qualitative evaluation shown in Figure 5, DashCLIP is able to create significantly more separation between the three classes compared to the baseline.
@@ 7.1. Pre-trained Embeddings:
Contextualized representations, or embeddings, derived from pre-trained models have demonstrated state-of-the-art performance across various natural language (Devlin, 2018) and vision tasks (Dosovitskiy, 2020). Early works such as Word2Vec (Mikolov, 2013) and Glove (Pennington et al., 2014) focused on learning word representations based on neighboring context and co-occurrence. ELMo (Sarzynska-Wawer et al., 2021) improved upon these with an LSTM-based architecture which enabled context-dependent word representations.
@@
The transformer architecture (Vaswani, 2017) gave rise to popular pre-trained language models like BERT (Devlin, 2018) and RoBERTa (Liu, 2019). Subsequent works leveraged this architecture to learn sentence-level embeddings (Reimers, 2019) (Gao et al., 2021). More recently, approaches have expanded to learn vision (Radford et al., 2021; Dosovitskiy, 2020) and multi-modal (Li et al., 2021, 2022, 2023) representations, enabling zero-shot transfer learning for vision-language tasks.
@@
Use of general datasets while training pre-trained models often leads to subpar performance on specific use cases, consequently requiring domain adaptation. In e-commerce and social media applications embeddings are widely used for personalized ranking and search. Nigam et al. (2019) learned semantic user representations using activity logs. MARN (Li et al., 2020) uses an adversarial network to generate multi-modal representations and achieved SoTA performance on several CTR prediction tasks. Baltescu et al. (2022) utilizes a combination of different engagement signals in a multi-task learning setup. Beyond these, numerous studies (Xv et al., 2022; Zhang et al., 2024; Pancha et al., 2022; Zhang et al., 2020) have tried to learn product and user representations on online platforms.
7.2. CTR Prediction:
@@
CTR prediction is crucial for e-commerce platforms and aims to predict the probability of a user clicking an impression. Earlier works relied on traditional machine learning methods including logistic regression (Richardson et al., 2007; Kumar et al., 2015), kernel SVM (Chang et al., 2010), Factorization Machines (FM) (Rendle, 2010; Juan et al., 2016; Pan et al., 2018), and gradient boosting (Ling et al., 2017) to predict CTR using different features. Deep Learning models lead to improved performance due to their ability to learn complex non-linear interactions: Product-based Neural Networks (PNN) (Qu et al., 2016) captured category interactions though embedding layers. The Wide and Deep architecture (Cheng et al., 2016) used a combination of wide cross-product transformations for feature interactions memorization and deep neural network for generalization, training both components jointly. Deep and Cross Network (DCM) (Wang et al., 2017) reduced the need for manual feature engineering by introducing the cross-network. DeepFM (Guo et al., 2017) combined DNNs with FMs to improve feature interaction. FiBiNET (Huang et al., 2019) used a bilinear function to dynamically learn feature interactions and demonstrates SoTA performance combining with both shallow and deep models. Finally, DCNV2 (Wang et al., 2021) improved on DCN, allowing large-scale learning in industrial settings. While many of these models emphasize architectural enhancements, our approach remains architecture-agnostic, focusing on generating semantic embeddings that can be seamlessly integrated into any architecture.